OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. (2x/y^2)^-4

OpenStudy (anonymous):

\[(\frac{2x}{y^2})^{-4}=(\frac{y^2}{2x})^4=\frac{y^8}{16x^4}\]